# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ Abstract
Over the past decade, deep learning has proven to be a highly effective tool for learning meaningful features from raw data. However, it remains an open question how deep networks perform hierarchical feature learning across layers. In this work, we attempt to unveil this mystery by investigating the structures of intermediate features. Motivated by our empirical findings that linear layers mimic the roles of deep layers in nonlinear networks for feature learning, we explore how deep linear networks transform input data into output by investigating the output (i.e., features) of each layer after training in the context of multi-class classification problems. Toward this goal, we first define metrics to measure within-class compression and between-class discrimination of intermediate features, respectively. Through theoretical analysis of these two metrics, we show that the evolution of features follows a simple and quantitative pattern from shallow to deep layers when the input data is nearly orthogonal and the network weights are minimum-norm, balanced, and approximate low-rank: Each layer of the linear network progressively compresses within-class features at a geometric rate and discriminates between-class features at a linear rate with respect to the number of layers that data have passed through. To the best of our knowledge, this is the first quantitative characterization of feature evolution in hierarchical representations of deep linear networks. Empirically, our extensive experiments not only validate our theoretical results numerically but also reveal a similar pattern in deep nonlinear networks which aligns well with recent empirical studies. Moreover, we demonstrate the practical implications of our results in transfer learning. Our code is available at https://github.com/Heimine/PNC_DLN.
Key words: deep representation learning, neural networks, intermediate features, feature compression and discrimination
@@ 1 Introduction
1.1 Background and Motivation
In the past decade, deep learning has exhibited remarkable success across a wide range of applications in engineering and science [74], such as computer vision [58, 108], natural language processing [107, 114], and health care [37], to name a few. It is commonly believed that one major factor contributing to the success of deep learning is its ability to perform hierarchical feature learning: Deep networks can leverage their hierarchical architectures to extract meaningful and informative features [^2] In deep networks, the (last-layer) feature typically refers to the output of the penultimate layer, which is also called representation in the literature. from raw data [10, 68]. Despite recent efforts to understand deep networks, the underlying mechanism of how deep networks perform hierarchical feature learning across layers still remains a mystery, even for classical supervised learning problems, such as multi-class classification. Gaining a deeper insight into this question will offer theoretical principles to guide the design of network architectures [53], shed light on generalization and transferability [79], and facilitate network training [118].
@@ 1.1 Background and Motivation
Towards opening the black box of deep networks, extensive empirical research has been conducted by investigating outputs at each layer of deep networks in recent years. An intriguing line of research has empirically investigated the role of different layers in feature learning; see, e.g., [1, 7, 25, 35, 83, 95, 126]. In general, these empirical studies demonstrate that the initial layers expand the intrinsic dimension of features to make them linearly separable, while the subsequent layers compress the features progressively; see Figure 1. For example, in image classification tasks, [1] observed that the features of intermediate layers are more and more linearly separable as we reach the deeper layers. Recent works [7, 95] studied the evolution of the intrinsic dimension of intermediate features across layers in trained networks using different metrics. They both demonstrated that the dimension of the intermediate features first blows up and later goes down from shallow to deep layers. More recently, [83] delved deeper into the role of different layers, and concluded that the initial layers create linearly separable intermediate features, and the subsequent layers compress these features progressively.
(a) A 7-layer MLP network with ReLU
(b) A 7-layer hybrid network
Figure 1: Illustration of numerical rank and training accuracy across layers. We train two networks with different architectures on the CIFAR-10 dataset: (a) A 7-layer multilayer perceptron (MLP) network with ReLU activation, (b) A hybrid network formed by 3-layer MLP with ReLU activation followed by a 4-layer linear network. For each figure, we plot the numerical rank of the features of each layer and training accuracy by applying linear probing to the output of each layer against the number of layers, respectively. The green shade indicates that the features at these layers are approximately linearly separable, as evidenced by the near-perfect accuracy achieved by a linear classifier. The definition of numerical rank and more experimental setups are deferred to Section 5.1.1.
@@
In the meanwhile, recent works [91, 52, 38, 127] provided more systematic studies into structures of intermediate features. They revealed a fascinating phenomenon during the terminal phase of training deep networks termed neural collapse (NC), across many different datasets and model architectures. Specifically, NC refers to a training phenomenon in which the last-layer features from the same class become nearly identical, while those from different classes become maximally linearly separable. In other words, deep networks learn within-class compressed and between-class discriminative features. Building upon these studies, a more recent line of work investigated the NC properties at each layer to understand how features are transformed from shallow to deep layers; see, e.g., [17, 53, 48, 43, 96]. In particular, [53] empirically showed that a progressive NC phenomenon governed by a law of data separation occurs from show to deep layers. The works [17, 43] provided empirical evidence confirming the occurrence of an NC property called nearest class-center separability in intermediate layers. [96] empirically showed that similar NC properties appear in intermediate layers during training, where the within-class variance decreases relative to the between-class variance as layers go deeper.
(a) A 7-layer nonlinear MLP
(b) A hybrid network with 3-layer MLP + 4-layer DLN
Figure 2: Visualization of feature compression & discrimination from shallow to deep layers. We consider the same setup as in Figure 1. For each network, we visualize the outputs of layers 1, 2, 4, and 6 on the CIFAR10 dataset using the 2-dimensional UMAP plot [81]. More experimental setups are deferred to Section 5.1.1.
In summary, extensive empirical results demonstrate that after feature expansion by initial layers, deep networks progressively compress features within the same class and discriminate features from different classes from shallow to deep layers; see also Figure 2. This characterization provides valuable insight into how deep networks transform data into output across layers in classification problems. Moreover, this insight sheds light on designing more advanced network architectures, developing more efficient training strategies, and achieving better interpretability. However, to the best of our knowledge, no theory has been established to explain this empirical observation of progressive feature compression and discrimination. In this work, we make the first attempt to provide a theoretical analysis of this observation based on deep linear networks (DLNs) to bridge the gap between the practice and theory of deep representation learning.
@@
Even though DLNs lack the complete expressive power of nonlinear networks, they possess comparable abilities for feature compression and discrimination to those found in the deeper layers of nonlinear networks, as indicated in Figures 1 and 2. By assessing the training accuracy and the numerical rank of intermediate features in both a nonlinear network and a hybrid network, [^3] For the hybrid network, we introduce nonlinearity in the first three layers and follow them with linear layers. it becomes apparent that the initial-layer features in both networks are almost linearly separable, evidenced by nearly perfect training accuracy achieved through linear probing. This phenomenon is further exemplified by the feature visualization presented in Figure 2. In the meanwhile, the linear layers in the hybrid network replicate the role of the counterpart in the nonlinear network to do feature compression and discrimination, as evidenced by the declining feature rank in Figure 1 and the increasing separation of different-class features in Figure 2 across layers in both types of networks.
Broadly speaking, DLNs have been recognized as valuable prototypes for investigating nonlinear networks due to the resemblance of certain behaviors in their nonlinear counterparts [1, 7, 83, 95] and their simplicity [3, 42, 99, 100]. For instance, [51] empirically demonstrated the presence of a low-rank bias at both initialization and after training for both linear and nonlinear networks. [100] showed that a DLN exhibits a striking hierarchical progressive differentiation of structures in its internal hidden representations, resembling patterns observed in their nonlinear counterparts.
@@
Although stacking linear layers in DLNs ultimately results in an end-to-end linear transformation from the input to the output, the overparameterization present in these DLNs sets them apart from a basic linear operator: increasing the depth of DLNs can significantly improve their generalization capabilities, enhance feature compression, and facilitate network training. Specifically, recent works demonstrated that linear over-parameterization by depth (i.e., expanding one linear layer into a composition of multiple linear layers) in deep nonlinear networks yields better generalization performance across different network architectures and datasets [40, 51, 72]. This is also corroborated by our experiments in Figure 3, where increasing the depth of linear layers of a hybrid network leads to improved test accuracy. Moreover, our results in Figure 12 imply that increasing depth of DLN also leads to improved feature compression. On the other hand, [3] showed that linear over-parameterization, such as increasing depth, acts as a preconditioner, which can accelerate convergence in DLNs. We refer interested readers to [88] for further discussion on the role of depth in DLNs.
@@
Figure 3: Depth of DLNs lead to better generalization performance. We train hybrid networks consisting of a 2-layer MLP with ReLU activation followed by $(L-2)$ linear layers on the FashionMNIST and CIFAR-10 datasets, respectively. As a reference, we also train nonlinear networks comprised exclusively of MLP layers. We plot the test accuracy against the different number of layers averaged over $5$ different runs. It is observed that adding linear layers and MLP layers can both yield better generalization performance. More experimental setups are deferred to Section 5.1.2.
@@ 1.2 Our Contributions
In this work, we study hierarchical representations of deep networks for multi-class classification problems. Toward this goal, we explore how a DLN transforms input data into a one-hot encoding label matrix from shallow to deep layers by investigating features at each layer. To characterize the structures of intermediate features, we define metrics to measure within-class feature compression and between-class feature discrimination at each layer, respectively. We establish a unified framework to analyze these metrics and reveal a simple and quantitative pattern in the evolution of features from shallow to deep layers:
@@
More specifically, we rigorously prove the above claim based upon the following assumptions:
  • Assumption on the training data. We study a $K$-class classification problem with balanced training data, where the number of training samples in each class is $n$. We assume that the training input samples are $\theta$-nearly orthonormal for a constant $\theta\in[0,1/4)$, and the feature dimension is larger than the total number of samples so that they are linearly separable.
  • Assumption on the trained weights. We assume that an $L$-layer DLN is trained such that its weights are minimum-norm, $\delta$-balanced, and $\varepsilon$-approximate low-rank due to the implicit bias of gradient descent, where $\delta,\varepsilon\in(0,1)$ is a constant.
We discuss the validity of our assumptions in Section 3.2. Based upon the above, we show in Section 3.1 that the ratio of within-class feature compression metric between the $(l+1)$-th layer and the $l$-th layer is determined by $O(\varepsilon^{2}/n^{1/L})$, implying a layerwise geometric decay from shallow to deep layers. On the other hand, we show that the metric of between-class discrimination increases linearly with respect to (w.r.t.) the number of layers, with a slope of $O((\theta+4\delta)/L)$. To the best of our knowledge, this is the first quantitative characterization of the feature evolution in hierarchical representations of DLNs in terms of feature compression and discrimination. In particular, this analysis framework can be extended to nonlinear networks. Finally, we substantiate our theoretical findings in Section 5 for both synthetic and real-world datasets, demonstrating that our claims are true for DLNs and also manifest in nonlinear networks empirically.
@@
In recent years, there has been a growing body of literature studying hierarchical feature learning to open the black box of deep networks. These studies include recent works on neural tangent kernel [60, 56], intermediate feature analysis [1, 83, 96], neural collapse [33, 53, 101, 110], and learning dynamic analysis [10, 12, 32], among others; we refer readers to Section 4 for a more comprehensive discussion. Our work contributes to this emerging area by showing that each layer of deep networks plays an equally important role in hierarchical feature learning, which compresses within-class features at a geometric rate and discriminates between-class features at a linear rate w.r.t. the number of layers. This provides a simple and precise characterization of how deep networks transform data hierarchically from shallow to deep layers. It also addresses an open question about neural collapse, and offers a new perspective justifying the importance of depth in feature learning. Moreover, our result explains why projection heads [21], which usually refer to one or several MLP layers added between the feature layer and final classifier during pretraining and discarded afterwards, can improve the performance of transfer learning on downstream tasks [79, 66, 44]. To sum up, our result provides important guiding principles for deep learning practices related to interpretation, architecture design, and transfer learning.
@@
Finally, we would like to highlight the differences and connections between our work and two closely related recent results [53, 100] in the following.
  • First, [53] empirically showed that a progressive NC phenomenon governed by a law of data separation occurs from shallow to deep layers. Specifically, they observed that in trained over-parameterized nonlinear networks for classification problems, a metric of data separation decays at a geometric rate w.r.t. the number of layers. This is similar to our studied within-class feature compression at a geometric rate, albeit with different metrics (see the remark after 1). However, they do not provide any theoretical explanation for the progressive NC phenomenon. While our research focuses on DLNs and may not provide a complete understanding of phenomena in nonlinear networks, our theoretical analysis of DLNs can offer insights into what we observe empirically in the deeper, nonlinear layers of such networks. This is supported by our findings presented in Figure 1, which show that linear layers can replicate the function of deep nonlinear layers in terms of feature learning.
  • Second, [100] reveals that during training, nonlinear neural networks exhibit a hierarchical progressive differentiation of structure in their internal representations—a phenomenon they refer to as progressive differentiation. While both their and our studies revealed and justified a progressive separation phenomenon based on DLNs, the focuses are entirely different yet highly complimentary to each other. Specifically, we investigate how a trained neural network separates data according to their class membership from shallow to deep layers after training, while they investigate how weights of a neural network change w.r.t. training time and their impact on class differentiation during training. This can be illustrated through the example in [100]. Suppose we train a neural network to classify eight items: sunfish, salmon, canary, robin, daisy, rose, oak, and pine. We study how these eight items are represented by the neural networks from shallow to deep layers in a trained neural network. In comparison, [100] explain why animals versus plants are first distinguished at the initial stage of training, then birds versus fish, then trees versus flowers, and then finally individual items in the whole training process.
1.3 Notation & Paper Organization

### v4

@@ Abstract
Over the past decade, deep learning has proven to be a highly effective tool for learning meaningful features from raw data. However, it remains an open question how deep networks perform hierarchical feature learning across layers. In this work, we attempt to unveil this mystery by investigating the structures of intermediate features. Motivated by our empirical findings that linear layers mimic the roles of deep layers in nonlinear networks for feature learning, we explore how deep linear networks transform input data into output by investigating the output (i.e., features) of each layer after training in the context of multi-class classification problems. Toward this goal, we first define metrics to measure within-class compression and between-class discrimination of intermediate features, respectively. Through theoretical analysis of these two metrics, we show that the evolution of features follows a simple and quantitative pattern from shallow to deep layers when the input data is nearly orthogonal and the network weights are minimum-norm, balanced, and approximately low-rank: each layer of the linear network progressively compresses within-class features at a geometric rate and discriminates between-class features at a linear rate with respect to the number of layers that data have passed through. To the best of our knowledge, this is the first quantitative characterization of feature evolution in hierarchical representations of deep linear networks. Moreover, our extensive experiments not only validate our theoretical results but also reveal a similar pattern in deep nonlinear networks, which aligns well with recent empirical studies. Finally, we demonstrate the practical value of our results in transfer learning. Our code is available at https://github.com/Heimine/PNC_DLN.
Keywords: deep representation learning, low-dimensional structures, linear discriminative representation, intermediate features, feature compression and discrimination
@@ 1 Introduction
In the past decade, deep learning has exhibited remarkable success across a wide range of applications in engineering and science (lecun2015deep), such as computer vision (he2016deep; simonyan2014very), natural language processing (sutskever2014sequence; vaswani2017attention), and health care (esteva2019guide), to name a few. It is commonly believed that one major factor contributing to the success of deep learning is its ability to perform hierarchical feature learning: deep networks can leverage their hierarchical architectures to extract meaningful and informative features [^1] In deep networks, the (last-layer) feature typically refers to the output of the penultimate layer, which is also called representation in the literature. from raw data (allen2023backward; krizhevsky2012imagenet). Despite recent efforts to understand deep networks, the underlying mechanism of how deep networks perform hierarchical feature learning across layers still remains a mystery, even for classical supervised learning problems, such as multi-class classification. Gaining deeper insight into this question will offer theoretical principles to guide the design of network architectures (he2023law), shed light on generalization and transferability (li2022principled), and facilitate network training (xie2022hidden).
@@
Towards opening the black box of deep networks, extensive empirical research has been conducted in recent years by investigating outputs at each layer of deep networks. An intriguing line of research has empirically investigated the role of different layers in feature learning; see, e.g., alain2016understanding; ansuini2019intrinsic; chen2022layer; masarczyk2023tunnel; recanatesi2019dimensionality; zhang2022all. In general, these empirical studies demonstrate that the initial layers expand the intrinsic dimension of features to make them linearly separable, while the subsequent layers compress the features progressively; see Figure 1. For example, in image classification tasks, alain2016understanding observed that the features of intermediate layers are increasingly linearly separable as we reach the deeper layers. Recent works (ansuini2019intrinsic; recanatesi2019dimensionality) studied the evolution of the intrinsic dimension of intermediate features across layers in trained networks using different metrics. They both demonstrated that the dimension of the intermediate features first blows up and subsequently goes down from shallow to deep layers. More recently, masarczyk2023tunnel delved deeper into the role of different layers and concluded that the initial layers create linearly separable intermediate features, while later layers compress these features progressively.
(a) A 8-layer MLP network with ReLU
(b) A 8-layer hybrid network
Figure 1: Illustration of numerical rank and training accuracy across layers. We train two networks with different architectures on the CIFAR-10 dataset: (a) A 8-layer multilayer perceptron (MLP) network with ReLU activation, (b) A hybrid network consisting of a 3-layer MLP with ReLU activation followed by a 5-layer linear network. For each figure, we plot the numerical rank of the features of each layer and the training accuracy obtained by applying linear probing to the output of each layer, both against the number of layers. The green shading indicates that the features at these layers are approximately linearly separable, as evidenced by the near-perfect accuracy achieved by a linear classifier. The definition of numerical rank and additional experimental details are deferred to Section 5.1.1.
@@
Meanwhile, recent works (papyan2020prevalence; Han2021; fang2021exploring; zhu2021geometric) have provided systematic studies on the structures of intermediate features. They revealed a fascinating phenomenon termed neural collapse (NC) during the terminal phase of training deep networks and across many different datasets and model architectures. Specifically, NC refers to a training phenomenon in which the last-layer features from the same class become nearly identical, while those from different classes become maximally linearly separable. In other words, deep networks learn within-class compressed and between-class discriminative features. Building upon these studies, a more recent line of work investigated the NC properties at each layer to understand how features are transformed from shallow to deep layers; see, e.g., ben2022nearest; he2023law; hui2022limitations; galanti2022implicit; rangamani2023feature. In particular, he2023law empirically showed that a progressive NC phenomenon, governed by a law of data separation, occurs from shallow to deep layers. rangamani2023feature empirically showed that similar NC properties emerge in intermediate layers during training, where within-class variance decreases relative to the between-class variance as layers go deeper.
(a) A 8-layer nonlinear MLP
(b) A hybrid network with 3-layer MLP + 5-layer DLN
Figure 2: Visualization of feature compression & discrimination from shallow to deep layers. We consider the same setup as in Figure 1. For each network, we visualize the outputs of layers 1, 2, 4, and 6 on the CIFAR-10 dataset using the 2-dimensional UMAP plot (mcinnes2018umap). Additional experimental details are deferred to Section 5.1.1.
In summary, extensive empirical results demonstrate that after feature expansion by initial layers, deep networks progressively compress features within the same class and discriminate features from different classes from shallow to deep layers; see Figure 2. This characterization provides valuable insight into how deep networks transform data into output across layers in classification tasks. Moreover, this insight sheds light on designing more advanced network architectures, developing more efficient training strategies, and achieving better interpretability. However, to the best of our knowledge, no theoretical framework has yet been established to explain this empirical observation of progressive feature compression and discrimination. In this work, we take a first step towards bridging this gap by providing a theoretical analysis based on deep linear networks (DLNs).
@@
Even though DLNs lack the strong expressive power of nonlinear networks, they possess comparable abilities for feature compression and discrimination to those observed in the deeper layers of nonlinear networks, as indicated in Figures 1 and 2. By evaluating the training accuracy and the numerical rank of intermediate features in both a nonlinear network and a hybrid network [^2] For the hybrid network, we introduce nonlinearity in the first few layers and follow them with linear layers., we observe that the initial-layer features in both networks are almost linearly separable, evidenced by nearly perfect training accuracy achieved through linear probing. This phenomenon is further illustrated by the feature visualization in Figure 2. Meanwhile, the linear layers in the hybrid network mimic the role of their counterpart in the nonlinear network by performing feature compression and discrimination, as evidenced by the decreasing feature rank in Figure 1 and the increasing separation of different-class features in Figure 2 across layers in both types of networks.
Broadly speaking, DLNs have been recognized as valuable prototypes for studying nonlinear networks, as they resemble certain behaviors of their nonlinear counterparts (alain2016understanding; ansuini2019intrinsic; masarczyk2023tunnel; recanatesi2019dimensionality) while maintaining simplicity (arora2018optimization; gidel2019implicit; saxe2019mathematical). For instance, huh2021low empirically demonstrated the presence of a low-rank bias at both initialization and after training for both linear and nonlinear networks. saxe2019mathematical showed that a DLN exhibits a striking hierarchical progressive differentiation of structures in its internal hidden representations, resembling patterns observed in their nonlinear counterparts.
@@
Although stacking linear layers in DLNs ultimately results in an end-to-end linear transformation from input to output, the overparameterization in these DLNs distinguishes them from a basic linear operator: increasing the depth of DLNs can significantly improve their generalization capabilities, enhance feature compression, and facilitate network training. Specifically, recent works have demonstrated that linear over-parameterization by depth (i.e., expanding one linear layer into a composition of multiple linear layers) in deep nonlinear networks yields better generalization performance across different network architectures and datasets (guo2020expandnets; huh2021low; kwon2023compressing). This is corroborated by our experiments in Figure 3, where increasing the depth of linear layers of a hybrid network leads to improved test accuracy. Moreover, our results in Figure 4 suggest that increasing the depth of DLN also leads to improved feature compression. We refer interested readers to nichani2021empirical for further discussion on the role of depth in DLNs.
@@
Figure 3: Depth of DLNs lead to better generalization performance. We train hybrid networks consisting of a 2-layer MLP with ReLU activation followed by (L2)(L-2) linear layers on the FashionMNIST and CIFAR-10 datasets, respectively. As a reference, we also train nonlinear networks comprised exclusively of MLP layers. We plot the test accuracy against the different number of layers averaged over 55 different runs. It is observed that adding either linear layers or MLP layers can improve generalization performance. More experimental details are deferred to Section 5.1.2.
(a) FashionMNIST
(b) CIFAR-10
Figure 4: Progressive feature compression on DLNs trained with default initialization and real datasets. Using the DLNs trained in Figure 3, we plot the within-class compression metrics $C_{l}$ (see Definition 1) against the layer indices. It is observed that progressive linear decay still (approximately) happens without the orthogonal initialization and datasets described in Assumption 2.
@@ 1.1 Our Contributions
In this work, we study hierarchical representations of deep networks for multi-class classification problems. Towards this goal, we explore how a DLN transforms input data into a one-hot encoding label matrix from shallow to deep layers by investigating features at each layer. To characterize the structures of intermediate features, we define metrics to measure within-class feature compression and between-class feature discrimination at each layer, respectively (see Definition 1). We establish a unified framework to analyze these metrics and reveal a simple and quantitative pattern in the evolution of features from shallow to deep layers:
@@
More specifically, we rigorously prove the above claim under the following assumptions:
  • Assumption on the training data. We study a $K$-class classification problem with balanced training data, where each class contains $n$ training samples. We assume that the training samples are $\theta$-nearly orthonormal for a constant $\theta\in[0,1/4)$, and the feature dimension is larger than the total number of samples so that they are linearly separable (see Assumption 2).
  • Assumption on the trained weights. We assume that an $L$-layer DLN is trained such that its weights are minimum-norm, $\delta$-balanced, and $\varepsilon$-approximately low-rank, due to the implicit bias of gradient descent, where $\delta,\varepsilon\in(0,1)$ are constants (see Assumption 3).
We discuss the validity of our assumptions in Section 3.2. Based upon these assumptions, we show in Theorem 4 that the ratio of the within-class feature compression metric between the $(l+1)$-th layer and the $l$-th layer is $O(\varepsilon^{2}/n^{1/L})$, implying a geometric decay from shallow to deep layers. Moreover, we show that the between-class discrimination metric increases linearly with respect to (w.r.t.) the number of layers, with a slope of $O((\theta+4\delta)/L)$. To the best of our knowledge, this is the first quantitative characterization of feature evolution in hierarchical representations of DLNs in terms of feature compression and discrimination. Notably, this analytical framework can be extended to nonlinear networks (jacot2025wide). Finally, we substantiate our theoretical findings in Section 5 on both synthetic and real-world datasets, demonstrating that our claims hold for DLNs and also manifest empirically in nonlinear networks.
@@
In recent years, there has been a growing body of literature studying hierarchical feature learning to open the black box of deep networks. These studies include works on neural tangent kernel (Jacot2018; huang2020dynamics), intermediate feature analysis (alain2016understanding; masarczyk2023tunnel; rangamani2023feature), neural collapse (dang2023neural; he2023law; Sukenik2023; tirer2022extended), and learning dynamic analysis (allen2023backward; bietti2022learning; Damian2022), among others. We refer readers to Section 4 for a more comprehensive discussion. Our work contributes to this emerging area by showing that each layer of deep networks plays an equally important role in hierarchical feature learning, which compresses within-class features at a geometric rate and discriminates between-class features at a linear rate w.r.t. the number of layers. This provides a simple and precise characterization of how deep networks transform data hierarchically from shallow to deep layers. It also addresses an open question about neural collapse and offers a new perspective justifying the importance of depth in feature learning; see the discussion in Section 3.1. Moreover, our result explains why projection heads (chen2020simple), which usually refer to one or several MLP layers added between the feature layer and final classifier during pretraining and discarded afterwards, can improve the performance of transfer learning on downstream tasks (li2022principled; kornblith2021why; galanti2022on). In summary, our result provides important guiding principles for deep learning practices related to interpretation, architecture design, and transfer learning.
@@
Finally, we highlight the differences and connections between our work and two closely related recent works (he2023law; saxe2019mathematical) as follows:
  • First, he2023law empirically showed that a progressive NC phenomenon governed by a law of data separation occurs from shallow to deep layers. Specifically, they observed that in trained over-parameterized nonlinear networks for classification problems, a metric of data separation decays at a geometric rate w.r.t. the number of layers. This is similar to our studied within-class feature compression at a geometric rate, albeit with different metrics (see the remark after Definition 1). However, they do not provide any theoretical explanation for the progressive NC phenomenon. While our research focuses on DLNs and may not provide a complete understanding of phenomena in nonlinear networks, our theoretical analysis offers insights into the empirical behavior of deeper nonlinear layers. This is supported by our findings presented in Figure 1, which demonstrate that linear layers can replicate the function of deep nonlinear layers in terms of feature learning.
  • Second, saxe2019mathematical reveals that during training, nonlinear neural networks exhibit a hierarchical progressive differentiation of structure in their internal representations—a phenomenon they refer to as progressive differentiation. While both their study and ours reveal and justify a progressive separation phenomenon based on DLNs, the focus of each study is entirely different yet highly complimentary to each other. Specifically, we investigate how a trained neural network separates data according to their class membership from shallow to deep layers after training, while they investigate how weights of a neural network change w.r.t. training time and their impact on class differentiation during training. This distinction can be illustrated by the example in saxe2019mathematical. Suppose a neural network is trained to classify eight items: sunfish, salmon, canary, robin, daisy, rose, oak, and pine. We study how these eight items are represented by the neural networks from shallow to deep layers in the trained neural network. In comparison, saxe2019mathematical explain why animals versus plants are first distinguished at the initial stage of training, then birds versus fish, then trees versus flowers, and finally individual items, throughout the training process.
1.2 Notation and Paper Organization